Calculate the volume in litres occupied by 2.12 mol
of nitric oxide (NO) at 6.54 atm and 76 0C
1
Expert's answer
2020-12-21T03:58:04-0500

p*V = n*R*T, where R is the gas constant

V = n*R*T/p = (2.12 mol)*(0.082 L*atm/mol/K)*(273.15+76 K)/(6.54 atm) = 9.28 L